OpenStudy (quickstudent):

How do I use the zero product property to solve this equation? -16t^2 + 28t + 8 = 0

OpenStudy (quickstudent):

@mathmate

OpenStudy (mathmate):

Zero product property means that if a*b = 0, then either a=0, b=0, or both equal to zero. When applied to factoring, say \(f(x)=x^2-3x+2=0\) Since \(f(x)=x^2-3x+2=(x-2)(x-1)\) which is a product of (x-2) and (x-1), then (x-2)=0, (x-1)=0, or both. This allows us to say x-2=0 means x=2, or x-1=0 means x=1. Therefore the solution to f(x)=0 is x=2 or x=1. You can now apply a similar procedure to the given problem in question.

OpenStudy (quickstudent):

So does that mean I have to factor the problem first to be able to find t?

OpenStudy (quickstudent):

@mathmate

OpenStudy (mathmate):

Yes, factoring would be the first step, once you know what to do afterwards.

OpenStudy (mathmate):

Do not forget to take out common factors of ALL the terms before attempting more elaborate factoring. Example: factor 2x^2-6x+4 = 2(x^2-3x+2)=2(x-2)(x-1)

OpenStudy (quickstudent):

I factored it to this now: -4(t - 2)(4t + 1) = 0 So, now I have to set each factor equal to 0, and solve for t, right?

OpenStudy (mathmate):

Exactly! (all factors except -4, of course!)

OpenStudy (quickstudent):

Yes, ok, then -4(t-2)=0 -4t-8=0 -4t-8+8=0+8 -4t=8 -4t/4=8/4 t=2 -4(2-2)=0

OpenStudy (quickstudent):

OR (4t + 1) = 0 4t + 1 = 0 4t + 1 - 1 = 0 - 1 4t = -1 4t/4 = -1/4 t = -1/4

OpenStudy (quickstudent):

Correct?

OpenStudy (mathmate):

Sorry I have ignored you for the last 8 minutes because I was helping someone else in the mean time! Perfect! Everything is correct and well done.

OpenStudy (quickstudent):

Thanks! :)

OpenStudy (mathmate):

You're welcome! :)